# Testing tidal theory using Gaia binaries: the red giant branch

### v1

@@ 6 Conclusions
Gaia DR3 (Gaia Collaboration et al., 2023) offers us an unprecedented opportunity to scrutinize tidal dissipation as stars evolve along the giant branch. We have made a direct comparison between observed giant binaries and the theoretical predictions, the latter obtained through a first-principled calculation of ‘f-mode’ tides, a non-zero frequency analogue to the well-known equilibrium tidal model. Fig. 11 summarizes our main results.
Overall, we find that f-mode tides, together with our adopted eddy viscosity, can circularize binaries out to a fixed ratio of $a/R\approx 4$, regardless of the evolutionary state. Any systems with dimensionless pericenters $\beta\equiv a(1-e)/R\leq 4$ should be effectively circular. In the meantime, Gaia giant binaries appear circularized out to $\beta\sim 6$ in the early RGB, and down to $\beta\sim 3$ in later stages. Moreover, like main-sequence binaries, Gaia binaries also exhibit ‘cool islands’, over-densities of circular orbits that extend well beyond the circularization limit. These islands reach $a/R\sim 10-15$. For the early RGB, this is about twice as far as predicted by our tidal calculations. Both of these features suggest, while f-mode tides can adequately describe the observational data during the late RGB, there needs to be an additional circularization process during the early RGB.
@@
We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC). This work has also made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC).
@@ Appendix A Equilibrium tides
$\omega^{2}\boldsymbol{\xi}+\rho_{0}^{-1}\nabla\cdot(2\rho_{0}\nu_{c}{\bf S})-\nabla(\boldsymbol{\mathcal{G}}\cdot\boldsymbol{\xi})-\nabla(c_{A}^{2}\nabla\cdot\boldsymbol{\xi})+\boldsymbol{\mathcal{A}}\nabla\cdot\boldsymbol{\xi}+\nabla\delta\Phi$ $=-\nabla U,$ (B1)
@@
Inserting potential terms $U_{mk}(r/R)^{2}Y_{2m}\exp[-{\rm i}\omega_{mk}t]$ from the series expansion of Equation C1, we compute dynamical ($\omega=\omega_{mk}\not=0$) solutions $\boldsymbol{\xi}_{mk}$ to Equations B1-B2 through an expansion in eigenmodes of the homogeneous problem (with $U=0$): $\boldsymbol{\xi}_{mk}=\sum_{\alpha}\tilde{a}_{\alpha}\boldsymbol{\xi}_{\alpha},$ where $\alpha$ labels oscillation modes with displacement eigenfunctions $\boldsymbol{\xi}_{\alpha}$ and natural frequencies $\omega_{\alpha}$. The amplitudes $\tilde{a}_{\alpha}(t)=a_{\alpha}\exp[-{\rm i}\omega_{mk}t]$ can then be determined approximately from (e.g., Schenk et al., 2001)
$a_{\alpha}\approx-(\omega_{\alpha}^{2}-\omega_{mk}^{2}-2\text{i}\gamma_{\alpha}\omega_{mk})^{-1}U_{mk}\int_{V^{*}}\rho_{0}\boldsymbol{\xi}_{\alpha}\cdot\nabla[(r/R)^{2}Y_{2m}]\text{d}V,$ (B3)
where $\gamma_{\alpha}$ gives the damping rate a given mode $\alpha$ would have if it were freely oscillating in isolation. Since many oscillations are driven simultaneously rather than in isolation, these ‘linear’ damping rates do not adequately describe viscous dissipation (Braviner & Ogilvie, 2015). However, their inclusion in Equation B3 regularizes resonances (where $\omega_{\alpha}=\omega_{mk}$). We then compute the total dissipation associated with tidal driving at the frequency $\omega_{mk}=k\Omega_{o}-m\Omega$ as
$\mathcal{D}(\omega_{mk})=\omega_{mk}^{2}\sum_{\alpha,\beta}\left(\frac{a_{\alpha}^{*}a_{\beta}}{\omega_{\alpha}\omega_{\beta}}\right)\int_{V^{*}}\rho_{0}\nu_{c}{\bf S}_{\alpha}^{*}:{\bf S}_{\beta}\text{d}V,$ (B4)
@@
We use a mode expansion that includes the degree $2$ f-mode and the ten lowest order p-modes. We compute these oscillations for each MESA timestep with GYRE (Townsend & Teitler, 2013). Because we focus on the later stages of the RGB, we place the inner boundary at the outer edge of the radiative core, and adopt GYRE’s ZERO_H inner boundary condition.
@@ Appendix C Tidal energy and angular momentum transfer
In the co-rotating, co-moving frame of a tidally perturbed star, the tidal potential imposed by a point mass, spin-aligned perturber can be written as
@@
$U_{nmk}=-\dfrac{GM^{\prime}}{a}\left(\frac{R}{a}\right)^{n}\frac{4\pi}{(2n+1)}Y_{n}^{m*}(\pi/2,0)X_{nmk}.$ (C2)
The degree $n=0$ part of a point-mass potential has been omitted from Equation (C1) because it does not produce a force, while the $n=1$ part is cancelled by fictitious forces in the co-moving Keplerian frame.
@@
$\delta\Phi=\sum_{\ell=0}^{\infty}\sum_{m=-\ell}^{\ell}\sum_{k=-\infty}^{\infty}\delta\Phi_{\ell m}^{k}\left(\frac{R}{r}\right)^{\ell+1}Y_{\ell}^{m}e^{-\text{i}\omega_{mk}t}.$ (C3)
@@
Here $\kappa_{mk}=\Im[k_{2m}^{2}]$ are the imaginary parts of $\ell=n=2$ tidal Love numbers. These can can be determined from Equations (B4) and (12) for a given $m$ and $k.$ Then, equating time derivatives of the orbital energy $E_{o}=-GMM^{\prime}/(2a)$, orbital angular momentum $L_{o}=-(1-e^{2})^{1/2}E_{o}/\Omega_{o}$, and stellar angular momentum $J$ with the rates of energy and angular momentum transfer ($\mathcal{P}$, $\mathcal{T}$, and $-\mathcal{T},$ respectively) leads to the evolutionary Equations (10)-(13).

### v2

@@ 6 Conclusions
Gaia DR3 (Gaia Collaboration et al., 2023) offers us an unprecedented opportunity to scrutinize tidal dissipation as stars evolve along the giant branch. We have made a direct comparison between observed giant binaries and theoretical predictions, the latter obtained through a first-principled calculation of ‘f-mode’ tides, a non-zero frequency analogue to the well-known equilibrium tidal model. Fig. 11 summarizes our main results.
Overall, we find that f-mode tides, together with our adopted eddy viscosity, can circularize binaries out to a fixed ratio of $a/R\approx 4$, regardless of the evolutionary state. Any systems with dimensionless pericenters $\beta\equiv a(1-e)/R\leq 4$ should be effectively circular. In the meantime, Gaia giant binaries appear circularized out to $\beta\sim 6$ in the early RGB, and down to $\beta\sim 3$ in later stages. Moreover, like main-sequence binaries, Gaia giant binaries also exhibit ‘cool islands’, over-densities of circular orbits that extend well beyond the circularization limit. These islands reach $a/R\sim 10-15$. For the early RGB, this is about twice as far as predicted by our tidal calculations. Both of these features suggest that, while f-mode tides can adequately describe the observational data during the late RGB, there needs to be an additional circularization process during the early RGB.
@@
[…]
We thank the anonymous reviewer for comments that improved the quality of the paper. We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), [funding reference number 513671]. This work has also made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC).
@@ Appendix A Equilibrium tides
$\frac{\partial^{2}\boldsymbol{\xi}}{\partial t^{2}}-\nabla(\boldsymbol{\mathcal{G}}\cdot\boldsymbol{\xi})-\nabla(c_{A}^{2}\nabla\cdot\boldsymbol{\xi})+\boldsymbol{\mathcal{A}}\nabla\cdot\boldsymbol{\xi}+\nabla\delta\Phi-\frac{1}{\rho_{0}}\nabla\cdot(2\rho_{0}\nu_{c}{\bf S})$ $=-\nabla U,$ (B1)
@@
Inserting potentials $U=\Re\{A(r/R)^{2}Y_{2m}\exp[-{\rm i}\omega_{mk}t]\}$ of nominal amplitude $A$, we compute dynamical ($\omega=\omega_{mk}\not=0$) solutions $\boldsymbol{\xi}_{mk}$ with the time dependence $\exp[-\text{i}\omega_{mk}t]$ to Equations B1-B2 through an expansion in eigenmodes of the homogeneous problem (with $U=0$): $\boldsymbol{\xi}_{mk}=\sum_{\alpha}\tilde{a}_{\alpha}\boldsymbol{\xi}_{\alpha},$ where $\alpha$ labels oscillation modes with displacement eigenfunctions $\boldsymbol{\xi}_{\alpha}$ and natural frequencies $\omega_{\alpha}$. The amplitudes $\tilde{a}_{\alpha}(t)=a_{\alpha}\exp[-{\rm i}\omega_{mk}t]$ can then be determined approximately from (e.g., Schenk et al., 2001)
$a_{\alpha}\approx-(\omega_{\alpha}^{2}-\omega_{mk}^{2}-2\text{i}\gamma_{\alpha}\omega_{mk})^{-1}AQ_{\alpha},$ (B3)
where $Q_{\alpha}=\int_{V}\rho_{0}\boldsymbol{\xi}_{\alpha}^{*}\cdot\nabla[(r/R)^{2}Y_{2m}]\text{d}V$, and $\gamma_{\alpha}$ gives the damping rate a given mode $\alpha$ would have if it were freely oscillating in isolation. Since many oscillations are driven simultaneously rather than in isolation, these ‘linear’ damping rates do not adequately describe viscous dissipation (Braviner & Ogilvie, 2015). However, their inclusion in Equation B3 regularizes resonances (where $\omega_{\alpha}=\omega_{mk}$). We then compute the amplitude-normalized dissipation associated with tidal driving at the frequency $\omega_{mk}=k\Omega_{o}-m\Omega$ as
$\mathcal{D}_{mk}=\frac{\omega_{mk}^{2}}{|A|^{2}}\sum_{\alpha,\beta}\left(\frac{a_{\alpha}^{*}a_{\beta}}{\omega_{\alpha}\omega_{\beta}}\right)\int_{V}\rho_{0}\nu_{c}{\bf S}_{\alpha}^{*}:{\bf S}_{\beta}\text{d}V,$ (B4)
@@
We use a mode expansion that includes the degree $2$ f-mode and the ten lowest order p-modes. We pre-compute these oscillations for each MESA timestep with GYRE (Townsend & Teitler, 2013). Because we focus on the later stages of the RGB, we place the inner boundary at the outer edge of the radiative core, and adopt GYRE’s ZERO_H inner boundary condition. We linearly interpolate the stellar profiles ($\rho_{0},$ $\ell_{c}$, $v_{c}$) and oscillation mode properties ($\omega_{\alpha}$, $\gamma_{\alpha},$ $Q_{\alpha}$, ${\bf S}_{\alpha}^{*}:{\bf S}_{\beta}$) required to compute tidal dissipation rates at each step of our integration of the secular tidal equations (10)-(13).
@@ Appendix C Tidal energy and angular momentum transfer
In the co-rotating, co-moving frame of a tidally perturbed star, the total tidal potential imposed by a point mass, spin-aligned perturber can be written as
@@
$U_{nmk}=-\dfrac{GM^{\prime}}{a}\left(\frac{R}{a}\right)^{n}\frac{4\pi}{(2n+1)}Y_{nm}^{*}(\pi/2,0)X_{nmk},$ (C2)
where $X_{nmk}$ are Hansen coefficients (see, e.g., eq. 15 in Zanazzi & Wu, 2021). The degree $n=0$ part of a point-mass potential has been omitted from Equation (C1) because it does not produce a force, while the $n=1$ part cancels in the frame co-moving with the primary.
@@
$\delta\Phi=\sum_{\ell=0}^{\infty}\sum_{m=-\ell}^{\ell}\sum_{k=-\infty}^{\infty}\delta\Phi_{\ell m}^{k}\left(\frac{R}{r}\right)^{\ell+1}Y_{\ell m}e^{-\text{i}\omega_{mk}t}.$ (C3)
@@
Here $\kappa_{mk}=\Im[k_{2m}^{2}]$ are the imaginary parts of $\ell=n=2$ tidal Love numbers. These can can be determined from Equations (B4) and (12) for a given $m$ and $k.$ Then, equating time derivatives of the orbital energy $E_{o}=-GMM^{\prime}/(2a)$, orbital angular momentum $L_{o}=-2(1-e^{2})^{1/2}E_{o}/\Omega_{o}$, and stellar angular momentum $J$ with the rates of energy and angular momentum transfer ($\mathcal{P}$, $-\mathcal{T}$, and $\mathcal{T},$ respectively) leads to the evolutionary Equations (10)-(13).